OpenStudy (anonymous):

What is √3 + 2√2 ÷ √3 − 2√2 ?

OpenStudy (anonymous):

Did you mean ? \[\frac{ \sqrt{3} + 2\sqrt{2} }{ \sqrt{3} - 2\sqrt{2} }\] ?

OpenStudy (anonymous):

Yes. :)

OpenStudy (anonymous):

In order to get rid off that ugly expression under the line, you need to multiply the whole fraction by \[\frac{ \sqrt{3} + 2\sqrt{2} }{ \sqrt{3} +2\sqrt{2} } \]

OpenStudy (anonymous):

they all cancel out and it is 1

OpenStudy (anonymous):

They dont cancel out and the result is not 1. Read it carefully, @Bhagyashree .